# Prompt-based Dynamic Token Pruning to Guide Transformer Attention in Efficient Segmentation

### v1

@@ 3 Experimental Results
Table 3: Ablation study, in terms of mean $DSC$ and standard deviation across all test samples, incorporating prompt-based adaptive pruning framework on different baseline models.
@@
Ventricle Myocardium Base Pruned
@@
Base Pruned Base Pruned Base Pruned
UNETR 0.552 $\pm$ 0.016 0.594 $\pm$ 0.001 0.695 $\pm$ 0.009 0.720 $\pm$ 0.017 0.777$\pm$ 0.006 0.782 $\pm$ 0.005 0.833 $\pm$ 0.009 0.875$\pm$ 0.013
Swin UNETR 0.64 $\pm$ 0.009 0.662 $\pm$ 0.018 0.794 $\pm$ 0.018 0.785 $\pm$ 0.006 0.846 $\pm$ 0.011 0.854 $\pm$ 0.017 0.885 $\pm$ 0.015 0.883 $\pm$ 0.007
TransUNET 0.640 $\pm$ 0.011 0.706 $\pm$ 0.007 0.769 $\pm$ 0.008 0.831 $\pm$ 0.005 0.871 $\pm$ 0.005 0.882 $\pm$ 0.009 0.824$\pm$ 0.011 0.825$\pm$ 0.018
SegFormer 0.684 $\pm$ 0.011 0.679 $\pm$ 0.013 0.822 $\pm$ 0.015 0.816 $\pm$ 0.011 0.898 $\pm$ 0.019 0.902 $\pm$ 0.004 0.812 $\pm$ 0.008 0.822 $\pm$ 0.002
Figure 1: Graphical plot, comparing (a) GFLOPs and (b) Token Sparsity of the baselines w.r.t. their pruned versions.
@@
Figure 3: (a) Sample input slice from ACDC, with token retention maps using (b) Swin UNETR, (c) SegFormer, (d) UNETR, and (e) TransUNET.
Row 1: ISIC, Row 2: ACDC, sample images. Red boxes denote comparison area.
Table 1 quantifies the experimental results of fixed threshold values (0.3, 0.5) and adaptive percentile-based threshold values ($25^{th},50^{th},75^{th}$) for $\tau$. Increasing the fixed value of $\tau$ from 0.3 to 0.5 reduces the computational cost. However, it may lead to over-pruning of relevant tokens due to a decrease in the mean DSC value. Lower percentile values ($25^{th}$ and $50^{th}$) retain comparatively more tokens than higher percentiles ($75^{th}$), as indicated by the increase in Giga Floating Point Operations (GFLOPs). $\tau$ at the $25^{th}$ percentile is found to have the optimal balance between accuracy and computational efficiency. Discarding 75% of the total tokens results in suboptimal segmentation performance due to information loss. Adaptive percentile-based thresholding provides greater flexibility to the token pruning mechanism than fixed threshold values, and achieves a superior balance between segmentation performance and computational costs.
Table 2 presents the results of the application of the proposed token pruning framework in multiple stages of UNETR. Pruning tokens at levels 1 and 2 drastically reduces the computational cost by 28%, with an associated increase in the mean value of DSC by 3%. Applying pruning to deeper levels of the network leads to a saturated segmentation performance without a significant decrease in computational efficiency. Therefore, early-stage pruning effectively removes irrelevant tokens and prevents them from spreading deeper into the network.
Table 3 summarizes the values of DSC for different baselines, as well as their pruned versions, in the ACDC and ISIC datasets. UNETR and TransUNET exhibit significant performance gains ($1\%-7\%$ approx.) in DSC values in both datasets. This suggests that the prompt-based spatial prior helps the model focus more on relevant regions. Swin UNETR and SegFormer demonstrate stable performance, even after substantial token removal, as seen in Fig. 1. Therefore, the proposed framework improves the overall segmentation accuracy by retaining relevant tokens for target structures.
Fig. 1(a) illustrates that the proposed framework significantly reduces computational costs across all models. The pruned variants of SegFormer and UNETR show the highest reduction in terms of GFLOPs. Fig. 1(b) depicts a significant decrease in the token density of the pruned variants, compared to their respective baselines. This elimination of irrelevant tokens leads to computationally efficient segmentation models. Fig. 2 demonstrates that the token relevance ranking distribution is naturally non-uniform, capturing differences in importance. The entropy-based weighting, on the other hand, is uniformly distributed. This is consistent with the theoretical analysis of the proposed approach, and confirms its efficacy in pruning.
Fig. 3 shows the token retention maps, generated for a sample input image, by the pruned version of the baselines. The variation in the percentage of pruned tokens for the same sample, across all models, highlights the specific feature learning mechanism of the different models. Swin UNETR, employing window-based self-attention, is found to retain the highest number of tokens. On the contrary, models utilizing full self-attention mechanisms like SegFormer exhibit comparatively higher token redundancy; as is evident from the lower percentage of tokens retained. The spread-out nature of the retained tokens implies that the model still retains global contextual information to some extent. This is necessary for efficient segmentation.
Fig. 4 qualitatively analyses the effectiveness of the proposed framework. The samples highlight cases where the pruned version effectively addresses issues of over- and under-segmentation in prediction by the baselines.
4 Conclusions and Discussion

### v2

@@ II-B Prompt-driven adaptive pruning
$\rho_{ik}>>\rho_{il}$ (6)
[…]
$r_{i}>>r_{j}$ (7)
Thus, the algorithm assigns higher scores to important tokens and lower scores to unimportant ones. Since, there exists a significant gap between the relevance scores, $\exists\>r_{min}=min_{i\in S_{T}}\{r_{i}\}$ and $r_{max}=max_{j\notin S_{T}}\{r_{j}\}$
By Eq. (7),
$r_{min}>r_{max}$
[…]
$\implies r_{min}>\sigma^{-1}(\tau)>r_{max}$
$\forall i\in S_{T}$, the relevance score $r_{i}>\sigma^{-1}(\tau)$ which ensures every token in $S_{T}$ is retained.
Thus, $S_{T}\subseteq S_{P}$. ∎
Corollary 1.
Given the token reduction factor $\Delta>0$, then $|S_{T}|<Z$
Proof.
By the fundamental property of typical sets we know,
$|S_{T}|\leq Z.2^{-\Delta}$ (8)
From the corollary we know,
$\Delta>0$ $\implies 2^{\Delta}>2^{0}$ (9)
$\implies 2^{\Delta}>1$ (10)
$\implies 2^{-\Delta}<1$ (11)
$\implies Z.2^{-\Delta}<Z$ (12)
[…]
$|S_{T}|\leq Z.2^{-\Delta}<Z\implies|S_{T}|<Z$ (13)
III Implementation details
This section summarizes the public datasets used in this study, the implementation details and a qualitative and quantitative analysis of experimental results in terms of performance metrics. Ablations and comparative study (with related state-of-the-art literature), demonstrated the efficacy of our PrATo framework.
III-A Datasets
The data set Automated Cardiac Diagnosis Challenge (ACDC) [28], acquired from the University Hospital of Dijon, France, contains 150 volumes of cardiac magnetic resonance images (MRI). Trained radiologists annotated the volumes for different cardiac components, viz. Right Ventricle (RV), Left Ventricle (LV), and Myocardium. The voxel intensities were clipped in the range [-170, 250] HU and subsequently normalized to the range of [0,1]. Data augmentation techniques, viz. random flipping, cropping and rotate were employed to expand the training dataset.
The data set ISIC [29] of the International Skin Imaging Collaboration is a collection of dermoscopic images for skin cancer research. Binary segmentation masks are created by experts where a pixel value of 0 indicates background while a pixel value of 1 denotes the lesion. The images were normalized, followed by the application of data augmentation techniques to enlarge the training data set.
III-B Performance Metrics and Loss function
The PrATo framework is evaluated on state-of-the-art ViT-based segmentation models, viz. UNETR, TransUNET, SegFormer and Swin UNETR. The models were trained in Python 3.9, using PyTorch and MONAI libraries, on a NVIDIA RTX A5500 GPU with 24GB of memory. The adopted train-test-validation split is 80:10:10. The framework uses the tightest possible bounding box to encompass the entire target region.
A combined loss function, incorporating Dice loss ($\mathbf{\mathcal{L}_{d}}$) with Categorical Cross-Entropy loss ($\mathbf{\mathcal{L}_{c}}$) [35] was used to train the models. The Dice loss function is commonly applied in image segmentation tasks that deal with class imbalance. Mathematically $\mathbf{\mathcal{L}_{d}}$ is expressed as
$\mathbf{\mathcal{L}_{d}}=\alpha-\sum_{\alpha=1}^{N}\left(\frac{2\sum_{i=1}^{V}\hat{y}_{\alpha,i}y_{\alpha,i}+\epsilon}{\sum_{i=1}^{V}\hat{y}_{\alpha,i}+y_{\alpha,i}+\epsilon}\right),$ (14)
where $N$ represents the total number of classes, with $\hat{y}_{\alpha,i}$ and $y_{\alpha,i}$ denoting the predicted and true values, respectively, for the $i$th voxel concerning class $\alpha$. $V$ signifies the total number of voxels in the input and $\epsilon$ is the additive smoothing parameter to prevent division errors by zero.
The categorical cross-entropy loss $\mathbf{\mathcal{L}_{c}}$ measures the difference between the probability distributions of the predicted and the ground truth map. It is expressed as
$\mathcal{L}_{c}=-\frac{1}{V}\sum_{i=1}^{V}\sum_{\alpha=1}^{C}y_{\alpha,i}\log{(\hat{y}_{\alpha,i})}.$ (15)
The composite loss function, using both Dice and categorical cross-entropy loss, leverages the respective benefits of each component. While $\mathbf{\mathcal{L}_{d}}$ mitigates the issue of class imbalance between foreground and background pixels, the $\mathcal{L}_{c}$ component adjusts the trade-off between False Positives ($FP$) and False Negatives ($FN$) on the predicted output map [35]. The composite loss function is
$\mathbf{\mathcal{L}(\{\hat{\gamma},\gamma\};\Phi})=\mathbf{\mathcal{L}_{d}(\{\hat{\gamma},\gamma\},\Phi})+\mathbf{\mathcal{L}_{c}(\{\hat{\gamma},\gamma\},\Phi}),$ (16)
where $\hat{\gamma}$ and $\gamma$ denote the predicted and the ground truth map, respectively. $\Phi$ represents the model parameters.
The Dice Score Coefficient ($DSC$), Intersection-Over-Union ($IoU$) and 95% Haursdorff distance ($HD95$) [36] were used to compare the segmentation performance between the original baselines and their adaptations after incorporating PrATo framework. The metrics are mathematically represented as
$DSC=\frac{2TP}{2TP+FP+FN},$ (17)
$IoU=\frac{TP}{TP+FP+FN},$ (18)
$HD95_{\alpha}(\hat{\gamma},\gamma)=\max_{\alpha\in C}\{\max_{\hat{y}_{\alpha,i\in\hat{\gamma}}}\min_{y_{\alpha,j\in\gamma}}\delta(\hat{y}_{\alpha,i\in\hat{\gamma}},y_{\alpha,j\in\gamma}),\\ \max_{y_{\alpha,i\in\gamma}}\min_{\hat{y}_{\alpha,j\in\hat{\gamma}}}\delta(\hat{y}_{\alpha,i\in\hat{\gamma}},y_{\alpha,j\in\gamma})\},$ (19)
where $TP$ represents True Positives and $\delta(.)$ corresponds to the Euclidean distance.
III-C Ablation Study
TABLE I: Component ablation study of PrATo on ACDC dataset. The best results are marked in bold.
PrATo Variants DSC mIoU mHD95
Ventricle Myocardium
Left Right
w/o Entropy 0.4923 0.6819 0.7586 0.5393 18.48
Naive Cropping 0.5272 0.7068 0.7739 0.5692 15.89
Proposed 0.6146 0.7171 0.819 0.6129 19.87
TABLE II: Ablation study of PrATo on ACDC dataset, over various threshold $\tau$.
$\mathbf{\tau}$ GFLOPs mDSC
Fixed 0.1 15.25 0.6131
0.2 15.78 0.6425
0.3 13.32 0.6701
0.5 9.07 0.6612
Percentile 25th 26.78 0.6933
50th 37.31 0.6952
75th 18.66 0.6575
TABLE III: Ablation study for spatial dimension of the RoI Align ($k$) on ACDC dataset.
$\mathbf{k}=$ 3 5 7 9
mDSC 0.6761 0.7169 0.6405 0.6627
The contribution of the core components of PrATo is presented in Table I. Two variants of PrATo are compared with the proposed framework, specifically, one without entropy-based weighting and the other replacing RoIAlign with Naive cropping. The second variant directly selects the tokens with center coordinates located within the box prompt. The proposed PrATo framework attains the highest $DSC$ value among all cardiac organs and demonstrates superior average $IoU$ relative to other variants. This indicates that the integration of RoIAlign for feature extraction and entropy-based token scoring effectively identifies relevant tokens. The variant lacking the entropy-weighting scheme demonstrates a notable decrease in $DSC$ and $mIoU$ values. This variant exclusively employs the unrefined features of RoIAlign to identify important tokens. Therefore, the entropy-weighting scheme is necessary to filter ambiguous features, leading to precise token selection. Naive cropping selects all tokens within the box indiscriminately, as it does not possess the advanced feature extraction capabilities of RoIAlign. This can be evidenced by the significant drop in $DSC$ values of the left ventricle and myocardium. Therefore, RoIAlign is the superior feature extraction approach for this task.
However, the average HD95 value ($mHD95$) is lower with the naive cropping approach than with the proposed framework. The fine-grained details especially around the edges are smoothed out by directly selecting all the underlying tokens within the box-prompt. This spatially averaged query directs the similarity mechanism to select contiguous groups of tokens, leading to predictions characterized by simpler boundaries. Despite the reduced penalties for smoother boundaries as indicated by the $HD95$ metric, the overall shape of the target structure is not accurately predicted, as demonstrated by the lower values of $DSC$ and $IoU$.
Table II quantifies the experimental results of fixed threshold values (0.1, 0.2, 0.3, 0.5) and adaptive percentile-based threshold values ($25^{th},50^{th},75^{th}$) for $\tau$. Lower values of $\tau$, viz. 0.1 and 0.2 exhibits a reduced value of mean DSC over the different classes in ACDC. This suggests that allowing low-relevance token processing tends to introduce noise, which impacts the final output quality. Although increasing the threshold value to 0.5 leads to a significant decrease in GFLOPs, it also results in a slight drop in segmentation accuracy, indicating over-pruning.
In contrast, percentile-based thresholding has higher segmentation accuracy, as evidenced by higher mDSC values. Lower percentile values ($25^{th}$ and $50^{th}$) retain comparatively more tokens than higher percentiles ($75^{th}$), as indicated by the increase in Giga Floating Point Operations (GFLOPs). $\tau$ at the $25^{th}$ percentile is found to have the optimal balance between accuracy and computational efficiency. Discarding 75% of the total tokens results in suboptimal segmentation performance due to information loss. Adaptive percentile-based thresholding provides greater flexibility to the token pruning mechanism than fixed threshold value and achieves a superior balance between segmentation performance and computational costs. The fixed threshold values remain constant for all inputs, making them inflexible to the distinct score distribution of each input. In contrast, percentile-based thresholding always preserves a constant proportion of relevant tokens in relation to the specific input. The risk of discarding essential tokens is minimal, thus maintaining the segmentation performance.
Table III presents the results of the ablation study performed to analyze the impact of different spatial dimensions of the region-specific characteristics derived from RoIAlign. Segmentation performance improves significantly from $k=3$ to 5. The $5\times 5$ feature map obtained by setting $k=5$ results in a high-dimensional query that allows the model to learn discriminative and fine-grained details from the bounded region of the box prompt. However, increasing the value of $k$ might include specific textures or noise unique to the small target region. Therefore, while finding its association with the rest of the tokens, the model fails to generalize and overfits to the specific input prompt. Consequently, the segmentation performance drops to higher values of $k$. Therefore, $k=5$ was chosen as the final value for the proposed framework.
IV Results and Discussion
TABLE IV: Comparative analysis of PrATo with other pruning frameworks on ACDC and ISIC datasets. The best results are marked in bold.
ACDC
DSC ISIC
Ventricle
Framework Left Right Myocardium mIoU mHD95 DSC IoU HD95 Inference time (ms)
DynamicViT 0.5115 0.5898 0.7302 0.4879 24.03 0.8493 0.7539 21.59 10.11
EvoViT 0.5418 0.681 0.8085 0.5561 24.5 0.8541 0.7566 8.30 59.15
DToP 0.4463 0.5598 0.6845 0.4447 21.7 0.8511 0.7449 35.41 15.23
Random Token Masking 0.5094 0.5927 0.7136 0.4905 18.85 0.8402 0.7377 13.83 2.66
STP 0.3958 0.4224 0.5677 0.3388 21.72 0.8544 0.7483 10.91 13.21
TRAM 0.4867 0.5448 0.6844 0.4562 20.08 0.6424 0.5514 40.02 3.07
PrATo 0.6146 0.7171 0.819 0.6129 19.87 0.8634 0.7678 17.34 3.55
Figure 3: Sample segmentation maps for comparing PrATo with other pruning frameworks on the ACDC dataset. (a) Input MRI image with overlay ground truth and sample outputs from (b) Dynamic ViT, (c) EvoViT, (d) STP, (e) DToP, (f) Random Token Masking, and (g) PrATo frameworks.
Figure 4: Sample segmentation maps for comparing PrATo with other pruning frameworks on the ISIC dataset. (a) Input dermoscopy image with overlay ground truth and sample outputs from (b) Random Token Masking, (c) Dynamic ViT, (d) EvoViT, (e) STP, (f) DToP, and (g) PrATo frameworks.
Table IV presents the experimental findings to understand the potential of the prompt-guided PrATo compared to other state-of-the-art pruning frameworks. The comparisons also include a crucial baseline, called Random Token Masking, which randomly drops a fixed number of tokens without any guidance from spatial priors like box-prompt. This highlights the impact of pruning without any auxiliary guidance, like spatial priors from box-prompts. Class-wise $DSC$ along with the average values $IoU$ and $HD95$ are reported for the ACDC dataset. PrATo shows superior performance in terms of $DSC$, achieving improvements of $7.28\%$, $3.61\%$ and $1.05\%$ for the left ventricle, right ventricle and myocardium, respectively, compared to the next best baseline. The mean $IoU$ exceeds that of the second-best method by nearly $7\%$. Random Token Masking demonstrates a marginal advantage over PrATo in terms of $mHD95$; however, its low $DSC$ and $mIoU$ values suggest under-segmentation of regions that align closely with the true organ regions, as depicted in Fig. 3(f).
A qualitative comparison is presented in Fig. 3 to visually assess segmentation maps from different frameworks. The output of PrATo is visually the closest to the ground truth compared to other approaches. The overall shapes of the target structures are accurately captured, indicating the importance of spatial priors in retaining relevant tokens. STP produces imprecise output with inaccurate boundaries and under-segmented regions of the left ventricle and myocardium, respectively, as seen in Fig. 3 (d). Dynamic ViT, EvoViT and DToP oversegments the left ventricle, as evident from the top row of Fig. 3 (b), (c), and (e).
PrATo achieves the best performance based on $DSC$ and $IoU$ scores on the ISIC dataset, as shown in Table IV. This indicates that the segmentation output of PrATO has the best overlap with the ground truth, accurately delineating the lesion region compared to other methods. However, EvoViT achieves the best $HD95$ score. This difference in boundary delineation efficacy between EvoViT and PrATo originates from their different pruning mechanisms. EvoViT retains irrelevant tokens to maintain the spatial grid structure, thus preserving finer boundary details. Although the box-prompt in PrATo serves as a robust prior to identifying the overall region, it represents a coarser form of guidance that may result in inaccuracies in boundary delineation.
Fig. 4 illustrates the sample segmentation maps of the PrATo and other pruning frameworks in the ISIC data set. The proposed framework generates visually precise segmentation output for both small and large lesions. The lesion boundaries are smooth and the overall shape aligns closely with the truth of the ground. Random Token Masking produces several under-segmented regions in the output. STP and DToP display numerous oversegmented areas characterized by irregular boundaries.
PrATo outperforms several high-accuracy methods in inference speed, such as EvoViT and Dynamic ViT, as detailed in Table IV. It is approximately 3x faster than Dynamic ViT and nearly 15x faster than the high-performing model EvoViT. The ability of PrATo to exhibit competitive performance with reduced inference times highlights the efficiency of the proposed prompt-guided mechanism. Although Random Token Masking and TRAM are comparatively faster than PrATo, their performance degradation makes them unsuitable for clinical use.
TABLE V: Quantitative analysis of the implementation of PrATo across various medical image segmentation baseline models on the ACDC and ISIC datasets. The DSC value is reported for each baseline and their corresponding pruned version.
@@ IV Results and Discussion
Ventricle Myocardium Baseline Pruned
@@
Baseline Pruned Baseline Pruned Baseline Pruned
UNETR 0.5521 0.5943 0.6952 0.7201 0.7773 0.7820 0.8330 0.8754
Swin UNETR 0.6400 0.6621 0.7940 0.7854 0.8462 0.8543 0.8850 0.8832
TransUNET 0.6401 0.7061 0.7693 0.8312 0.8710 0.8821 0.8242 0.8253
SegFormer 0.6841 0.6791 0.8221 0.8163 0.8981 0.9024 0.8121 0.8220
Figure 5: Graphical plot, comparing (a) GFLOPs and (b) Token Sparsity of the segmentation models with respect to their pruned versions.
PrATo was integrated with various ViT-based state-of-the-art segmentation models to demonstrate its generalizability and model-agnostic nature. Table V summarizes the values of DSC for the segmentation models and their pruned versions using PrATo in the datasets ACDC and ISIC. The proposed framework is applied to the output of ViT blocks for non-hierarchical models such as UNETR and TransUNET and subsequently to the patch merging stage for hierarchical models like Swin UNETR and SegFormer. This ensures that the architectural operations in the models are not compromised and the reduced set of tokens is propagated to subsequent stages. UNETR and TransUNET exhibit significant performance gains ($1\%-7\%$ approx.) in DSC values in both datasets. This suggests that the prompt-based spatial prior helps the model focus more on relevant regions. Swin UNETR and SegFormer demonstrate stable performance, even after substantial token removal, as seen in Fig. 5. Therefore, the proposed framework improves the overall segmentation accuracy by retaining relevant tokens for target structures.
Models with a full self-attention mechanism, such as UNETR and TransUNET, evaluate the association between all possible pairs of tokens. However, this might include processing irrelevant tokens. The proposed pruning framework functions as a regularizer, enabling the model to focus on relevant tokens corresponding to the target structures. This improves the segmentation performance. In contrast, models with local self-attention, like Swin UNETR and SegFormer, are optimized to attend the relevant local features, leading to efficient segmentation performance. Therefore, the primary benefit of PrATo incorporated variants of such models is the computational savings rather than performance improvement.
Fig. 5(a) illustrates that the proposed framework significantly reduces computational costs across all models. The pruned variants of SegFormer and UNETR show the highest reduction in terms of GFLOPs. Fig. 5(b) depicts a significant decrease in the token density of the pruned variants, compared to their respective baselines. This elimination of irrelevant tokens leads to computationally efficient segmentation models. Fig. 6 qualitatively analyzes the effectiveness of the proposed framework. The samples highlight cases where the pruned version effectively addresses issues of over- and under-segmentation in prediction by the baselines. The experiments suggest that PrATo is a versatile framework to improve the efficacy of different segmentation models.
Row 1: ISIC, Row 2: ACDC, sample images. Red boxes denote the comparison area.
Figure 7: Sample token retention maps of PrATo from ACDC (a)-(b) and ISIC (c)-(d) datasets. The red box highlights the target region.
@@
Fig. 7 visualizes the sample token retention maps from both ACDC and ISIC datasets to gain insight into the effectiveness of the proposed approach. The retained tokens (marked by the yellow patches) correspond to the relevant spatial regions associated with the target structures. PrATo retains a dense set of tokens for the ACDC dataset, as shown in Fig. 7(b), which is highly concentrated within the target region. This preserves the finer details related to the different cardiac organs. In contrast, the framework intelligently adapts to the homogeneous texture of the skin lesions and prunes redundant tokens related to the uniform structure of the region of interest (Fig. 7(d)). This illustrates that PrATo seamlessly adapts its pruning mechanism based on the characteristics of the target structures. Fig. 8 visualizes the effectiveness of the scoring and weighting mechanism of PrATo. Fig. 8(a) illustrates that the token relevance ranking distribution is non-uniform, with a small subset of highly relevant tokens. The flat “ True Uniform ” line marked in red represents a vast majority of tokens have relevance scores close to zero. Therefore, these scores might lead to instability while generating the pruning mask. Fig. 8(b) demonstrates the weight distribution after the proposed entropy-based weighting scheme. The transformed weights follow the true uniform line, which highlights that the proposed method effectively transforms the skewed weights to a nearly uniform distribution. Therefore, such a stable weighting system indicates a robust pruning process as it prevents a few outlier tokens with extreme scores from skewing the thresholding process.
Figure 9: Qualitative results of PrATo under different prompting conditions. (a) Oversized, (b) tight-boxed, (c) partial and (d) misleading prompts.
TABLE VI: Quantitative analysis of PrATo under different prompting conditions on the ACDC dataset. The best results are marked in bold.
Approach mDSC mIoU mHD95
Tight Prompts 0.7169 0.6129 19.87
Oversized Prompts 0.6571 0.552 21.43
Partial Prompts 0.675 0.5763 17.16
Misleading Prompts 0.6528 0.5489 19.31
The results of the robustness analysis of the proposed framework under erroneous prompting conditions are presented in Table VI. The performance of PrATo was analyzed across varying scenarios, viz., tight, oversized, partial and misleading (incorrectly placed box) prompts. The quantitative results suggest that PrATo exhibits graceful degradation when the box prompts are placed incorrectly. The partial prompts show significant robustness with minimal drop in the mean values of $DSC$ and $IoU$. This is because partial prompts still provide high-quality features related to a portion of the target structure. The global context learned from the self-attention mechanism of ViT helps to accurately predict the remaining regions. Oversized prompts include background information along with the target region, resulting in over-segmentation, as evidenced in Fig. 9(a). The misleading prompts guide the attention of the framework away from the target region, thereby leading to the lowest average values of $DSC$ and $IoU$. However, the segmentation model learns robust feature representations, strong enough to override the misleading spatial priors. This leads to a reasonable performance. Although there are several segmentation errors(Fig. 9(d)), the overall structure is largely preserved. Thus, PrATo can handle imperfect prompts common in interactive segmentation without collapsing the overall segmentation performance of the underlying segmentation model.
V Conclusion